OpenStudy (anonymous):

At breakfast, Gino drank 2/3 of the orange juice in a container. He drank 15 ounces of orange juice. A. Write a multiplication equation that you can use to find the amount of orange juice that was in the container before Gino had breakfast. HELP and explain how to do it

OpenStudy (anonymous):

@Data_LG2

OpenStudy (anonymous):

PLZ help

OpenStudy (anonymous):

hold on ill c wat i can do to help

OpenStudy (anonymous):

OJ(2/3) = 15 OJ = 15 / (2/3) OJ = 15 * 3/2 OJ = 27.5 oz All integers can be expressed as rational numbers but not all rational numbers (exp π/2, 4/7, etc) can be expressed as integers. However since there exists an infinite number of both, the question, " Are there more rational numbers than integers?" is meaningless. It's like asking, "Are there more unicorns than married bachelors?" Now having said that, your math teacher may fallaciously try to argue that since integers (I) are a subset of rational numbers it is somehow "self evident" that the set "R" (rational numbers) is numerically the larger of the two sets. The key here is to recognize that a logical relationship implied in integers being a subset of rational numbers does not necessarily lead one to conclude that a numerical relationship of the form (R > I) therefore must follow. We argue that since both sets are equal (i.e infinite) than the actual numerical relation between both sets would be R = I. Words like "more" and "subset" often come cognitively "pre-loaded" suggesting superior/subordinate numerical relationships that may not, under scrutiny, actually exist.

OpenStudy (anonymous):

~please post your sources. thank you.

OpenStudy (anonymous):

I don't have answer choices for this

OpenStudy (anonymous):

im sorry @Data_LG2

OpenStudy (anonymous):

no worries (:

OpenStudy (anonymous):

@torifamelol give me a sec, im trying to understand it

OpenStudy (anonymous):

let x be the amount of orange juice that was in the container before Gino had breakfast. so you'll have \(\sf \frac{2}{3} x = 15\ ounces \), can you solve for x? i'm just doing trial and error for now.

OpenStudy (anonymous):

2/3= 3/2=1.5

OpenStudy (anonymous):

hmm.. 3/2=1.5 and 1.5x15=22.5 like bribro said. Is that correct?

OpenStudy (anonymous):

i pretty much explained it and @Data_LG2 made it easier all you gotta do is x and find the answer

OpenStudy (anonymous):

I got this THANKS!:D I have another one I am confused on though.

OpenStudy (anonymous):

Thank u guys

OpenStudy (anonymous):

your welcome anytime